Post date:
Monday, April 16, 2012 - 11:43
Description:
The Internet traffic will be moved to the backup firewall to validate a new firewall firmware release.
Users are invited to report anomalies with connections to external sites to neteng@cern.ch.
All the TCP connections to external sites will be reset at the time of the moves.
Reason for this intervention:
Validation of new firmware with production traffic.
Effective from date:
Tuesday, April 17, 2012 - 07:30
End date:
Tuesday, April 24, 2012 - 07:30
Service Element Affected:
Network Infrastructure Service
Impact on service:
Intervention should be transparent
Posted by:
IT-CS
Unit responsible for resolution:
IT Department